### v2

# Agentar-DeepFinance-100K: A Large-Scale Financial Dataset via Systematic Chain-of-Thought Synthesis Optimization

Abstract
Recent advancements in large language models (LLMs) have demonstrated remarkable general reasoning capabilities, holding significant potential for applications in the financial domain, a field that requires robust and reliable reasoning. It has been demonstrated that distilling high-quality chain-of-thought (CoT) rationales from advanced general reasoning models offers a promising and efficient path to the financial reasoning model. However, existing CoT synthesis methods suffer from shallow CoT sampling, leaving the question of how to construct a well-designed knowledge space for finance reasoning unexplored. In this paper, we present Agentar-DeepFinance-100K , a large-scale financial reasoning dataset characterized by its systematic CoT synthesis optimization. We first introduce a comprehensive CoT synthesis pipeline featuring Multi-perspective Knowledge Extraction (MKE) and Self-Corrective Rewriting (SCR) to generate exhaustive and deep financial reasoning trajectories. Furthermore, a systematic investigation, termed CoT Cube, is conducted to analyze critical factors that influence CoT effectiveness, such as necessity, length and synthesizer, yielding valuable insights for high-quality financial CoT construction. Experiments demonstrate that models trained on our Agentar-DeepFinance-100K achieve significant improvements on financial benchmarks. We publicly release Agentar-DeepFinance-100K , hoping to advance the research in financial reasoning models. The project page is available at https://github.com/antgroup/Agentar-DeepFinance-100K.
1 Introduction
Figure 1: Overview of our proposed CoT synthesis pipeline (left) and systematic investigation of factors that impact CoT effectiveness (right).
Pre-trained on extensive corpora of world knowledge, large language models (LLMs) have showcased impressive generalization capabilities, marking a significant milestone on the way to artificial general intelligence (AGI). When further fine-tuned on domain-specific data, financial LLMs [37, 33, 40] also exhibit substantial potential in tasks such as financial sentiment analysis [1], financial time series [41] and quant investment [14]. Despite these advancements, such models still face significant challenges in solving financial problems that demand complex reasoning. Encouragingly, recent progress in general reasoning models, such as OpenAI o1 [20], DeepSeek-R1 [8] and QwQ [39], has pushed beyond the limitations of traditional LLMs by explicitly modeling chain-of-thoughts (CoTs). These developments suggest a promising and efficient approach for building financial LLMs, i.e., distilling high-quality financial CoTs from general reasoning models.
Recent works [16, 47] have demonstrated that synthesizing financial CoTs from general reasoning models can significantly enhance financial reasoning capabilities. However, as illustrated in Figure 1 (A), existing methods for CoT synthesis primarily involve directly sampling CoTs from reasoning models for selected question-answer (QA) pairs, followed by rejection sampling for quality filtering. These approaches often result in shallow financial reasoning knowledge extraction, lacking the comprehensive and challenging knowledge required for robust financial reasoning. Additionally, synthesized CoTs are rarely subjected to systematic examinations to identify critical factors (e.g., length) that influence their effectiveness in reasoning enhancement. Without such investigation, it remains unclear how CoTs can be optimized to effectively drive improvements in financial reasoning.
To this end, this paper introduces a comprehensive pipeline for synthesizing CoTs, as shown in Figure 1 (B). Beginning with the curated seed corpora, our methodology integrates the Multi-perspective Knowledge Extraction (MKE) strategy and the Self-Corrective Rewriting (SCR) mechanism for CoT sampling. On the one hand, this pipeline demonstrates enhanced capabilities in extracting exhaustive domain knowledge for financial reasoning. On the other hand, it enables the sampling of more complex and challenging questions, leading to more intricate reasoning trajectories. Furthermore, a systematic investigation is conducted regarding critical factors influencing CoT effectiveness, such as necessity, length and synthesizer. As shown in Figure 1, we named it CoT Cube. Empirical evaluations of these components provide valuable insights into CoT-driven optimization of financial reasoning models and offer empirical guidelines for constructing high-performance training datasets.
Employing this method, we present Agentar-DeepFinance-100K , a large-scale financial dataset designed to push the boundary of financial reasoning models. Beyond the synthesizing CoTs, Agentar-DeepFinance-100K also incorporates multi-dimensional metadata annotation, such as complexity, quality and task type. Agentar-DeepFinance-100K is distinguished from existing financial reasoning datasets by its systematic CoT synthesis optimization. This includes building a comprehensive knowledge space for financial reasoning through our pipeline and refining CoT generation through the exploration of CoT Cube. Furthermore, to bridge the disparity between training data and real-world financial interactions, we additionally curated a dataset annotated by financial experts to reflect the financial capabilities required in real-world scenarios.
The remainder of this report is organized as follows. Section 2 provides an overview of our proposed Agentar-DeepFinance-100K . Section 3 details our comprehensive CoT synthesis pipeline. Section 4 presents an in-depth investigation into the factors influencing the effectiveness of CoT and the ablation study for the CoT synthesis pipeline. Section 5 discusses related works with this report. Section 6 summarizes the technical contributions and future works.
2 Dataset Overview
Figure 2: Illustration of the format of Agentar-DeepFinance-100K , which comprises three components: (1) the question, (2) the solution, including CoT and the final answer and (3) metadata, which encompasses multi-dimensional annotations such as complexity, quality and language.
Dataset Format. As illustrated in Figure 2, each sample in the dataset is accompanied by rich metadata in addition to the question and solution. This metadata is annotated using LLMs through a carefully designed manual prompt. Annotating this metadata not only provides a comprehensive understanding of the dataset but also can serve as a basis for subsequent experiments. Specifically, the metadata encompasses the following dimensions:
  • Content: A domain-specific classification of financial topics (e.g., stock).
  • Ability: Classification of the sample’s capability, including language, reasoning, knowledge, mathematics, code, instruction following and agents.
  • Complexity: Numerical score (1-10) represents the difficulty by considering the knowledge depth, multi-step reasoning requirements and instruction compliance challenges.
  • Quality: Score (1-10) evaluating response accuracy, completeness and procedural clarity against financial domain standards. Samples with a quality rating below 8 are excluded from the dataset.
  • Language: Identification of the language used in the sample, such as Chinese (zh) and English (en).
  • Task: A hierarchical decomposition of the task (e.g., Text Creation $\rightarrow$ Marketing Copy Generation).
Dataset $N_{\text{Q}}$ $N_{\text{R}}$ $N_{\text{A}}$Samples Proportion
Open-source Dataset
FinCorpus 86.25 1775.09 330.61 61,574 $53.64\%$
FinCUGE 99.17 498.44 232.40 18,527 $16.14\%$
Finance-Instruct-500K 151.49 920.85 558.15 10,343 $9.01\%$
FinQA 1051.05 1928.63 189.91 5,012 $4.37\%$
FinancialData 14.85 896.85 803.57 4,047 $3.53\%$
Quant-Trading-Instruct 65.97 2665.40 903.90 89 $0.08\%$
In-house Dataset
FinCAS 69.58 315.24 1.03 15,190 $13.23\%$
FinRA - - - - -
Table 1: Overview of Agentar-DeepFinance-100K data sources.
Data Source. The Agentar-DeepFinance-100K dataset is constructed from two primary sources: 99K samples from open-source data and 16K samples from our in-house data. For the open-source portion, extensive QA pairs were gathered from publicly available datasets, including FinCorpus [31], Finance-Instruct-500K [6], FinCUGE [17], FinQA [2], FinancialData [4] and Quant-Trading-Instruct [18]. These datasets underwent rigorous filtering to ensure quality and relevance to financial contexts, and instances overlapping with test sets in the benchmarks were systematically removed. The in-house data consists of two primary components: financial real annotations (FinRA) and content related to financial compliance and security (FinCAS). Table 1 summarizes the data sources, presenting some statistics such as average question length $N_{\text{Q}}$, reasoning trajectory length $N_{\text{R}}$ and answer length $N_{\text{A}}$. Reasoning trajectories were generated using QwQ-Plus.
Figure 3: Task composition.
Figure 4: Complexity distribution.
Task Diversity Analysis. Figure 4 presents the task composition within our Agentar-DeepFinance-100K , which encompasses six primary task domains: Knowledge QA, Natural Language Processing (NLP), Analysis & Interpretation, Math, Compliance & Security and Text Generation. The task distribution exhibits a significant concentration, with Knowledge QA and NLP collectively accounting for over $70\%$ of the dataset. In contrast, Text Generation tasks represent the smallest proportion, accounting for less than $1\%$. This distribution pattern aligns with observed user query tendencies in the financial domain, where interactions predominantly involve straightforward queries related to basic financial knowledge and information.
Distribution of Complexity. Figure 4 illustrates the distribution of complexity scores assigned to the Agentar-DeepFinance-100K , with scores ranging from 1 to 10. These complexity scores are generated by a large language model using a carefully crafted prompt designed to capture the relative complexity in solving the problems, rather than directly correlating with the difficulty of the tasks themselves. The results indicate that the complexity scores within the dataset are primarily concentrated in the range of 6 to 7, corresponding to medium-level questions. Additionally, a significant proportion of questions are scored between 3 and 5, which reflects lower levels of complexity. Furthermore, a smaller subset of the dataset includes questions with complexity scores exceeding 7, indicative of higher levels of complexity.
3 Dataset Construction
3.1 Overview
Our proposed dataset construction pipeline is illustrated in Figure 5. The pipeline consists of three main stages: (1) Starting from seed corpora, multi-perspective knowledge extraction is first adopted to generate diverse QA pairs. (2) Sample candidate CoTs from current advanced large reasoning models (LRMs) [8, 20, 39] and perform answer consistency verification against golden answers. (3) Implement self-corrective rewriting on QA pairs that cannot be sampled correctly. Our comprehensive pipeline facilitates systematically producing high-quality, diverse and challenging reasoning trajectories.
Figure 5: Illustration of the pipeline for constructing Agentar-DeepFinance-100K .
3.2 Seed Corpora
The seed corpora serve as the foundational component for dataset construction, requiring extensive coverage across diverse financial domains. Existing public financial datasets [2, 3, 16] provide structured knowledge from standardized materials (e.g., examinations), restricting their capacity to capture real-world user interactions. To bridge this gap, we systematically curated a large proprietary dataset with real domain expert annotations.
3.3 Multi-perspective Knowledge Extraction
The knowledge extraction phase aims to generate high-quality and rich QA pairs, which can model both explicit and implicit knowledge from the seed corpus. To achieve this, we propose the Multi-perspective Knowledge Extraction (MKE), comprising three extraction approaches:
  • Q2A (Direct Curation): It serves as the foundation by directly harvesting well-structured QA pairs from the seed corpora. Subsequently, removing duplicates and filtering out low-quality content ensures the relevance and quality.
  • A2Q (Counterfactual Augmentation): The A2Q mechanism implements counterfactual exploration to enhance the knowledge space’s density and causal connectivity. By perturbing original answers through semantic negation and contextual antonym substitution, adversarial answer variants are constructed. Subsequently, we adopt LRMs to generate questions for these answers, followed by a multi-stage verification protocol assessing semantic coherence, logical consistency and answer faithfulness. This bidirectional adversarial synthesis expands the neighborhood of core knowledge points.
  • T2Q (CoT Knowledge Mining): During the thinking phase of LRM, additional knowledge points may be introduced in CoTs. Sometimes, these knowledge points are not common sense but crucial to the reasoning process. Therefore, we developed a methodology to extract this latent knowledge from CoTs. The approach involves summarizing CoTs, prompting LRMs to identify key processes, and creating QA pairs. This process captures emergent knowledge points that bridge the gap between surface-level QA and deep reasoning requirements.
Our multi-perspective knowledge extraction framework achieves exhaustive and comprehensive knowledge modeling: Q2A preserves canonical patterns, A2Q strengthens causal connectivity through counterfactual perturbations, and T2Q uncovers implicit reasoning dependencies. The experiment in Section 4.4 demonstrates the effectiveness of this paradigm.
3.4 CoT Sampling and Verification
After multi-perspective knowledge extraction, numerous QA pairs are generated. To further construct high-quality reasoning traces, we first sample multiple CoTs and corresponding answers for each QA pair. Noticing that regular expression-based verification methods are insufficient for financial QA due to answer format variability (e.g., monetary expressions), we employ a lightweight model to match the answers. Since the verification task is simple and well-defined, a small model is sufficient and more cost-effective than a larger one. Only rigorously verified QA pairs with matching numerical precision and logical coherence are retained, ensuring the dataset’s reliability for complex financial reasoning tasks.
3.5 Self-Corrective Rewriting
While modern LRMs exhibit impressive performance, significant challenges persist in addressing complex financial problems. A common practice in existing methodologies is to discard QA pairs that lack correctly sampled answers, thereby ensuring the overall accuracy of the dataset. However, these discarded QA pairs often involve intricate reasoning processes, which could serve as valuable resources for enhancing model capabilities. Therefore, we came up with this problem: Can flawed reasoning trajectories be transformed into effective learning opportunities through appropriate correction mechanisms?
To accomplish this, we introduce Self-Corrective Rewriting (SCR) to enable the model to refine its answers by giving it some insights from the golden answers. More specifically, the SCR comprises two main steps: (1) Reflection phase: The model examines the disparities between its incorrect answer $A$ and the golden answer $A^{*}$ to generate diagnostic reflections CoT $T_{rf}$, which contains potential mistakes. (2) Rewriting phase: The reflection CoT $T_{rf}$ is merged with the original reasoning trace $T$. The model then continues the generation process based on this merged CoT to produce a new CoT $T_{rw}$ and the revised answer $A_{rw}$. Subsequently, $A_{rw}$ is verified against the correct answer $A^{*}$. If the verification is successful, exit the SCR process. The final CoT output consists of the original CoT, with the reflection CoT $T_{rf}$ and rewritten CoT $T_{rw}$ from each iteration alternately appended. The ultimate answer output is $A_{rw}$. If the verification fails, the correction cycle persists until the iteration limit is reached.
After SCR, more challenging questions can be correctly sampled. It is observed that the CoTs generated by SCR become much longer than before. Experiments in Section 4.4 have shown that integrating these challenging questions’ CoTs obtained through SCR can significantly boost the model’s reasoning capabilities.
4 Experiment
Evaluation Setup. For rapid experiments, we randomly selected 60K samples from our rigorously quality-filtered dataset. The experiments were conducted on the ms-swift framework [46], with the Qwen2.5-7B-Instruct model trained for 3 epochs. Model performance was assessed on three financial benchmarks: FinQA [2], Fin-Eva [29] and a self-sampled in-domain subset. Among these, Fin-Eva consists of multiple-choice and true/false questions that can be verified through rule-based evaluation. In contrast, FinQA and our in-domain test set include open-ended questions with answers that are difficult to parse using regular expressions, necessitating the use of the LLM-as-a-judge evaluation approach. Fin-Eva’s test set is divided into two parts: the Ant Group and the SUFE data sources. The results presented in this paper are based on data from the Ant Group source. Additionally, to assess general reasoning capabilities, we evaluated two widely-used benchmarks: MATH-500 [10] and GPQA-Diamond [25]. The pass@1 accuracy is reported with $1$ and $8$ responses per query for MATH-500 and GPQA-Diamond, respectively.
4.1 Investigation on CoT necessity
Existing works typically synthesize CoTs exclusively for questions with formally verifiable answers, leaving it unclear which financial question types can benefit from CoTs. In this section, we conduct experiments identifying financial questions where CoT yields gains. Specifically, we explore two classification schemes: one based on task type and the other on question difficulty. Our experiments are performed on the rigorously filtered subset of 60K samples.
Method FinQA Fin-Eva
Qwen2.5-7B-Instruct 63.47 83.70
Simple $w/o$ CoT 67.65 84.21
$-w/\text{-}$ CoT 68.96 85.01
Hard $w/o$ CoT 63.03 75.80
$-w/\text{-}$ CoT 69.31 85.96
Table 2: CoT necessity analysis on difficulty.
Difficulty. To validate the relationship between question difficulty and the necessity of CoTs, we classify questions into simple and hard subsets based on correctness labels from superior LLMs (e.g., Qwen2.5-14B-Instruct). For a fair comparison, equal sample sizes were maintained for both subsets. Table 2 presents the results with and without CoTs on these two types of questions. It demonstrates that CoT yields gains on both simple and hard questions and improves more significantly on hard problems. Notably, when fine-tuning on hard questions without CoTs, a performance decline was observed. It suggests that complex financial reasoning may not be sufficiently learned by models lacking explicit CoT guidance.
CoT Text Gen. Analysis & Interp. NLP Compliance & Sec. Math Knowledge QA Average
$w/o$ 74.55 81.36 76.71 91.25 31.71 78.32 72.32
$-w/\text{-}$ 83.64 88.05 77.21 92.02 61.15 79.01 80.18
Table 3: CoT necessity analysis on task type.
Task Type. We partition the dataset into 6 subsets aligned with the financial task taxonomy defined in Section 2. For each subset, experiments with and without CoT integration are conducted, reserving $10\%$ for testing. Additionally, the security compliance test set also includes the security compliance cases from Fin-Eva for more objective evaluation. As shown in Table 3, CoT integration consistently improves performance across all tasks. Notably, reasoning-intensive tasks, such as financial mathematics, exhibit substantial gains, demonstrating the critical role of CoT in complex financial reasoning.
Takeaway 4.1 for CoT Necessity Incorporating CoT consistently improves model performance regardless of task type and difficulty, particularly in complex reasoning tasks (e.g., math) and difficult problems.
4.2 Investigation on CoT Synthesizer
With the rapid advancement of large reasoning models, the research community has witnessed the emergence of numerous models, characterized by enhanced general reasoning capabilities. Many existing studies [16, 47] predominantly employ DeepSeek-R1 for distilling financial CoTs, with limited investigation into CoT synthesizers. This lack of exploration leaves unanswered questions regarding the impact of different reasoning models on CoT synthesis and which model is best suited for synthesizing financial CoTs.
Model Dense Fin-Eva FinQA General
Acc Res. Length Acc Res. Length MATH GPQA
DeepSeek-R1 91.42 770.14 75.41 675.46 98.20 61.62
QwQ-Plus 90.99 627.43 72.97 1831.49 97.20 54.55
Qwen3-235B-A22B 90.01 428.54 74.54 1302.86 94.60 62.12
Qwen2.5-7B-Instruct 83.70 1.00 63.47 239.97 73.20 37.88
Teacher Model #Samples/Avg. Length Fin-Eva FinQA General
Acc Res. Length Acc Res. Length MATH GPQA
DeepSeek-R1 21,040/1564.63 85.32 754.54 66.52 1270.96 80.00 41.41
QwQ-Plus 19,323/1795.99 85.53 702.41 68.79 2196.81 81.60 42.42
Qwen3-235B-A22B 20,322/1522.92 85.45 640.81 66.43 1469.85 80.00 42.42
Table 4: Performance comparison of various CoT synthesizers, including their intrinsic performance and the performance of corresponding distilled student models. Additionally, the average response lengths on Fin-Eva and FinQA are reported.
In this section, we employ multiple general reasoning models to synthesize CoTs for the same financial dataset, aiming to identify the most effective synthesizer. Specifically, a subset of 25K samples was randomly selected, with each model generating four CoTs and corresponding answers per question. Only questions containing the correct sampled answer were retained for analysis. Table 4 summarizes the experimental results, comparing both the intrinsic reasoning capabilities of the models and their effectiveness in distilling knowledge to improve student models. It reveals a key insight: a good model is not necessarily good at teaching students.
It shows that DeepSeek-R1 demonstrated superior performance not only in financial reasoning tasks but also in general knowledge evaluation, compared with QwQ-Plus and Qwen3-235B-A22B. Moreover, DeepSeek-R1 sampled more questions correctly during the process of sampling CoTs for the 25K samples. It can also be observed that DeepSeek-R1 produced CoTs in shorter and more stable lengths when answering questions in Fin-Eva and FinQA. However, the performance of smaller models fine-tuned using CoTs generated by DeepSeek-R1 was suboptimal. Interestingly, QwQ-plus, a model with relatively weaker financial and general capabilities, achieved the best overall performance. Specifically, compared with DeepSeek-R1, the model distilled from QwQ-Plus achieved improvements of +0.2 and +2.2 on Fin-Eva and FinQA, respectively, as well as increases of +1.6 and +1.0 on MATH and GPQA. Further analysis revealed that CoTs generated by QwQ-Plus tend to be longer than those synthesized by DeepSeek-R1. This characteristic likely contributed to the superior performance of models fine-tuned with QwQ-Plus-generated CoTs. Consequently, we adopt QwQ as the primary generator for financial CoT synthesis.
Takeaway 4.2 for CoT Synthesizer The effectiveness of a reasoning model as a CoT synthesizer does not always align with its intrinsic reasoning performance.
4.3 Investigation on CoT Length
To investigate the impact of CoT length, we employ a prompt-based approach to generate CoTs of varying lengths from the same CoT synthesizer. Specifically, except for the original CoTs, which are recognized as long CoTs, we generate two additional shorter versions of CoTs: (1) Middle-length CoTs: Inspired by CCoT [26], we append the phrase "Be concise" to the instruction, which significantly reduces the CoT length. (2) Short-length CoTs: We select samples where the model exhibits overthinking and manually simplify them. The model then generates shorter CoTs via in-context learning using these simplified samples.
Based on the findings in Section 4.2, the CoT synthesizer employed in this section is QwQ-Plus. The experimental results presented in Table 5 validate the effectiveness of the prompt-based approach in controlling CoT length while maintaining overall performance. In terms of CoT length, the generated responses on FinQA were reduced by $25.0\%$ and $29.6\%$ compared to the long CoTs. The only notable difference was that using short CoTs led to a $1.5$ decrease in accuracy on Fin-Eva.
CoT Length Fin-Eva FinQA
Acc Res. Length Acc Res. Length
Long 90.99 627.43 72.97 1831.49
Medium 91.07 562.98 72.45 1374.59
Short 89.51 459.48 72.97 1288.94
Table 5: Performance comparison of QwQ-Plus under different CoT lengths on Fin-Eva and FinQA. The medium- and short-length CoTs are generated via our prompt-based CoT length control method.
CoT Length Avg. Length Fin-Eva FinQA
Acc Res. Length Acc Res. Length
Long 1352.16 85.53 702.41 68.79 2196.81
Medium 1248.42 85.40 742.90 65.13 1698.63
Short 1020.26 85.19 574.00 65.56 1587.76
Table 6: Comparison of the models distilled from QwQ-Plus with different CoT lengths. The average CoT lengths in the training corpora are reported.
Table 6 illustrates the performance of the model distilled using CoTs of three different lengths. It further confirms the effectiveness of our prompt-based CoT length control method, with medium-length and short-length CoTs resulting in reductions of $7.7\%$ and $24.5\%$, respectively. Notably, the model trained with the longest CoT demonstrates superior performance on both the Fin-Eva and FinQA benchmarks. Specifically, it achieves improvements of +$3.7$ and +$3.2$ on FinQA compared to models trained with medium-length and short CoTs, respectively. In conclusion, while reducing the CoT length in the training corpus can result in more concise model responses, it may also compromise the model’s overall performance.
Takeaway 4.3 for CoT Length Distilling CoTs with reduced length allows the trained model to provide more concise responses, but may also hurt performance. Financial reasoning requires long CoTs.
4.4 Ablation Study
Figure 6: Ablation Study for Multi-perspective Knowledge Extraction.
Figure 7: Ablation Study for Self-Corrective Rewriting.
Impact of Multi-perspective Knowledge Extraction. To evaluate the effectiveness of our proposed Multi-perspective Knowledge Extraction (MKE), we conducted experiments with the 60K samples subset (denoted as Q2A). In this process, we employed the A2Q and T2Q synthesis methods within MKE to create two new datasets: one with 19K A2Q samples and another with 34K T2Q samples. The experimental results are shown in Figure 7. It shows that training with only the Q2A dataset yields a $+3.09$ improvement on Fin-Eva. Moreover, incorporating the T2Q dataset with Q2A enhances performance by $+1.12$, and incorporating A2Q with Q2A yields $+1.47$. Furthermore, using the A2Q and T2Q datasets for joint training yields a significant improvement of +$1.92$.
Impact of Self-Corrective Rewriting. To evaluate the effectiveness of SCR, we conducted experiments using the FinQA training set (5K samples) and reported the results on the FinQA test set. We sampled 4 candidate answers for each question through QwQ-Plus and found that 1,337 samples lacked correct answers after verification. Figure 7 presents our experimental results. For these 1,337 incorrectly sampled cases, we compared training performance using self-consistency and random sampling. The results show improved performance when training on incorrectly sampled data, with random sampling outperforming self-consistency. In addition, after training with 862 samples containing correct answers generated after applying SCR, the performance continued to improve by $+1.74$, validating the efficacy of SCR.
5 Related Work
Reasoning Model. The initial developments in reasoning models stemmed from approaches involving few-shot prompting with decomposed reasoning steps [34] and zero-shot activation through trigger phrases like "Let’s think step by step" [13]. Progress in this field was further driven by RL-based train-time scaling [42] and search-based test-time scaling [28], indicating a promising direction for enhancing the reasoning capacities of LLMs. A significant milestone in this direction is OpenAI o1 series [20], which showcased impressive reasoning capabilities. Subsequently, a series of open-source o1 reproductions emerged, such as OpenR [32], LLaVA-o1 [38], o1-Coder [44], LLaMA-Berry [43] and Journey Learning [24, 11, 12]. Recently, DeepSeek-R1 [8] introduced long CoT cold start and rule-based RL, proving that simply verifying the final answer can allow the model to automatically explore and generate long CoTs, thereby improving its reasoning ability. Subsequent research endeavors have focused on replicating R1 (e.g., OpenR1 [5]) or applying it to various domains, such as R1-Omni [45] in sentiment analysis, MedVLM-R1 [21] in medical imaging and Fin-R1 [16] in finance.
Financial LLMs. The Finance LLM is designed to enhance the general LLM’s capabilities in the financial domain. Early efforts, such as FinBERT [1], built on BERT and focused on tasks like sentiment analysis and financial text classification. With the success of GPT models, domain-specific models such as BloombergGPT [35], FinMA [37], FinGPT [33], InvestLM [40] have emerged, leveraging instruction tuning and curated datasets to improve performance on financial tasks. Recent advancements have integrated reasoning capability, as seen in O1-based models such as XuanYuan-FinX1-Preview [30] and Fino1 [23], which demonstrate enhanced proficiency in handling complex financial reasoning tasks. Furthermore, Fin-R1 [16] and DianJin-R1 [47] distill cold-start CoTs from DeepSeek-R1 [8] and perform rule-based RL on hard-to-sample examples.
CoT Synthesis. Initially, the CoTs generated by the few-shot [34] and zero-shot [13] methods are short. Moreover, the CoT examples adopted in few-shot methods are manually designed. Subsequent advancements [22, 7] integrated reward models (e.g., PRM [15]) to guide the search of high-quality CoTs. More recently, with the emergence of powerful reasoning models [8, 20, 39], distillation [9, 19, 5] from these models emerged as a cost-effective strategy for CoT Synthesis. This process involves querying large language models (LLMs) to generate CoTs and answers, followed by verification to ensure consistency. Additionally, post-processing techniques, such as token compression for efficiency [36] and length extension for richer reasoning [27], have further refined the distilled CoTs.
6 Conclusion
This paper presents Agentar-DeepFinance-100K , a large-scale financial reasoning dataset constructed with a systematically optimized chain-of-thought (CoT) synthesis framework. We propose a pipeline incorporating multi-perspective knowledge extraction and self-corrective rewriting to produce comprehensive reasoning trajectories. Additionally, we systematically investigate critical factors influencing CoT effectiveness, providing empirical insights for high-quality financial CoT construction. Experiments demonstrate the effectiveness of our Agentar-DeepFinance-100K , with significant improvements on financial reasoning benchmarks of models trained on our dataset. Future work will focus on exploring multi-modal financial reasoning, incorporating visual financial content such as charts.
Appendix
Figure 8: An example for T2Q (CoT knowledge) in multi-perspective knowledge extraction.
Figure 9: An example for A2Q (Counterfactual Augmentation) in multi-perspective knowledge extraction.
Figure 10: Examples for our in-house financial dataset FinRA.